i need to know the answer to this multi stepped equation:
-10=w-32+6

Answer by elima(1433)

-10=w-32+6
First you want to add -32+6; but remember 32 is a negative number so you actually are going to subtract; 32-6=26. Now is the 26 negative or positive? look at the original equqtion, which number is the larger - 32, is it negative or positive? It is negative so the 26 is negatve, -26.
Now you have:
-10=w-26
Now you want to get rid of the -26 so you get the w by itself, so you will add it to both sides.
26+-10=w-26+26
16=w